# Multimodal Evaluation of Russian-language Architectures

### v2

@@ G.1 Model Baselines Details
Qwen/Qwen2-VL-72B-Instruct 0.54 0.49 / 0.61 0.57 / 0.69 0.34 / 0.55
Qwen/Qwen2.5-VL-72B-Instruct 0.506 0.57 / 0.49 0.65 / 0.58 0.53 / 0.21
Qwen/Qwen3-VL-8B-Instruct 0.483 0.53 / 0.51 0.61 / 0.61 0.46 / 0.18
Qwen/Qwen2.5-VL-7B-Instruct 0.474 0.49 / 0.52 0.58 / 0.63 0.41 / 0.23
Qwen/Qwen2.5-Omni-7B 0.42 0.42 / 0.52 0.44 / 0.59 0.24 / 0.32
Qwen/Qwen2.5-VL-3B-Instruct 0.418 0.39 / 0.48 0.47 / 0.58 0.25 / 0.34
Qwen/Qwen3-VL-2B-Instruct 0.396 0.41 / 0.46 0.51 / 0.57 0.2 / 0.24
openbmb/MiniCPM-o-2_6 0.341 0.33 / 0.44 0.41 / 0.53 0.13 / 0.19
Qwen/Qwen2.5-Omni-3B 0.332 0.3 / 0.45 0.33 / 0.53 0.12 / 0.26
Qwen/Qwen2-VL-7B-Instruct 0.298 0.08 / 0.5 0.12 / 0.61 0.04 / 0.43
OpenGVLab/InternVL3-9B-Instruct 0.293 0.26 / 0.28 0.25 / 0.29 0.33 / 0.35
OpenGVLab/InternVL3-9B 0.276 0.28 / 0.27 0.26 / 0.26 0.32 / 0.27
Qwen/Qwen2-VL-2B-Instruct 0.198 0.06 / 0.37 0.06 / 0.47 0.05 / 0.18
OpenGVLab/InternVL3_5-4B 0.188 0.27 / 0.14 0.28 / 0.14 0.24 / 0.06
OpenGVLab/InternVL3-2B 0.134 0.09 / 0.2 0.11 / 0.24 0.07 / 0.09
OpenGVLab/InternVL3_5-2B-Instruct 0.133 0.14 / 0.14 0.15 / 0.19 0.1 / 0.09
llava-hf/LLaVA-NeXT-Video-7B-hf 0.122 0.09 / 0.2 0.09 / 0.23 0.03 / 0.08
@@
G.2 Human Baseline Details

### v3

@@ F.1 Model Baselines Details
Qwen2.5-VL-72B-Instruct 0.63 0.57 / 0.64 0.65 / 0.73 0.53 / 0.63
Qwen3-VL-8B-Instruct 0.58 0.53 / 0.60 0.61 / 0.69 0.46 / 0.56
Qwen2-VL-72B-Instruct 0.56 0.49 / 0.62 0.57 / 0.70 0.34 / 0.62
Qwen2.5-VL-7B-Instruct 0.52 0.49 / 0.56 0.58 / 0.67 0.41 / 0.44
Qwen2.5-Omni-7B 0.44 0.42 / 0.55 0.44 / 0.61 0.24 / 0.39
Qwen2.5-VL-3B-Instruct 0.43 0.39 / 0.49 0.47 / 0.60 0.25 / 0.37
Qwen3-VL-2B-Instruct 0.42 0.41 / 0.48 0.51 / 0.59 0.20 / 0.31
Qwen3-Omni-30B-A3B-Instruct 0.41 0.00 / 0.00 0.64 / 0.72 0.48 / 0.63
MiniCPM-o-2_6 0.37 0.33 / 0.49 0.41 / 0.57 0.13 / 0.30
Qwen2.5-Omni-3B 0.34 0.30 / 0.46 0.33 / 0.54 0.12 / 0.27
InternVL3-9B-Instruct 0.32 0.27 / 0.30 0.25 / 0.32 0.33 / 0.44
InternVL3-9B 0.32 0.28 / 0.31 0.26 / 0.31 0.32 / 0.42
Qwen2-VL-7B-Instruct 0.30 0.09 / 0.50 0.12 / 0.61 0.04 / 0.45
InternVL3_5-4B 0.29 0.27 / 0.31 0.28 / 0.32 0.24 / 0.32
LLaVA-NeXT-Video-7B-hf 0.13 0.09 / 0.21 0.09 / 0.24 0.03 / 0.09
@@
F.2 Human Baseline Details